emma has 50 coins in her pocket and their value is $6.30. If emma only has nickels and quarters, how many of each does she have?
Found 2 solutions by Rongtian, TimothyLamb:
Answer by Rongtian(7)

x=quarter amount
y=nickel amount
(-5)(x+y)=10(-5)
25x+5y=225
20x=175
no solution

Answer by TimothyLamb(4379)

x = nickels
y = quarters
---
x + y = 50
5x + 25y = 630
---
put the system of linear equations into standard form
---
x + y = 50
5x + 25y = 630
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = nickels = 31
y = quarters = 19
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php